Fin the absolute value of 4|2y-7|+5=9
Found 2 solutions by Alan3354, ptth176:
Answer by Alan3354(69443)
Answer by ptth176(50)

4|2y-7|+5=9
=>|2y-7|=1
=>2y-7=1 or 2y-7=-1
=>y=4 or y=3